# Performance measurements of the electromagnetic calorimeter and readout electronics system for the DarkQuest experiment

### v1

@@ 7 Background Rates in DarkQuest’s Experimental Cavern
We applied a low energy threshold of 50 MeV to allow for a complete analysis of the event rates throughout the energy spectrum. The resulting distribution showed little variation between channels and a total event rate of roughly $4\cdot 10^{3}$ counts per second, as shown in Fig. 11. We note that 15 of 16 channels were functional during these studies due to physical damage to a cable connection for one of the channels. This event rate, however, was mostly comprised of low energy events. Fig. 11 also shows the event rate as a function of energy threshold and demonstrates that the vast majority of these events are less than 1 GeV. This implies that we can increase the energy threshold required for each event and significantly reduce the event rate. We found that the average number of events per spill triggering on an energy greater than 5 GeV is 0.343 and above 10 GeV is 0.015. Although this represents only 4 modules of the 648 total for the EMCal, a naive scaling would lead to an expected background rate of around 55 events per spill above 5 GeV and less than 3 events per spill above 10 GeV for the full-sized EMCal. This is three orders of magnitude below the limitations of the current SpinQuest DAQ system and provides evidence that there are not unanticipated backgrounds (e.g. neutrons) and that the trigger rate can be kept to a very low level.
Figure 11: Left: Total number of counts above an energy value for all 4 modules averaged over all spills. The beam intensity was $10^{12}$ for the duration of the run. Right: Event rate over time for the first 30 minutes of this run, clearly showing intensity spikes from the recurrent beam spills.
@@ 8 Conclusions
The commissioning of the DarkQuest EMCal was performed with test beam data. The studies on linearity, resolution, and response to various particles confirm the suitability of the EMCal and its newly developed readout electronics for the DarkQuest experiment. Light attenuation with neutral density filters and electrical attenuation were found to give similar performance. Background rates in the real experiment were also studied and were found to be within the constraints of the existing readout system.

### v2

@@ 7 Background Rates in DarkQuest’s Experimental Cavern
We applied a low energy threshold of 50 MeV to allow for a complete analysis of the event rates throughout the energy spectrum. The resulting distribution showed little variation between channels and a total event rate of roughly $4\cdot 10^{3}$ counts per second, as shown in Fig. 12. We note that 15 of 16 channels were functional during these studies due to physical damage to a cable connection for one of the channels. This event rate, however, was mostly comprised of low energy events. Fig. 12 also shows the event rate as a function of energy threshold and demonstrates that the vast majority of these events are less than 1 GeV. This implies that we can increase the energy threshold required for each event and significantly reduce the event rate. We found that the average number of events per spill triggering on an energy greater than 5 GeV is 0.343 and above 10 GeV is 0.015. Although this represents only 4 modules of the 648 total for the EMCal, a naive scaling would lead to an expected background rate of around 55 events per spill above 5 GeV and less than 3 events per spill above 10 GeV for the full-sized EMCal. This is three orders of magnitude below the limitations of the current SpinQuest DAQ system. While the SpinQuest background is not uniform, as this naive scaling would suggest, the extremely low background estimate provides evidence that there are not unanticipated backgrounds (e.g. neutrons) and that the trigger rate can be kept to a sufficiently low level.
Figure 12: Left: Total number of counts as a function of the threshold on energy for all 4 modules averaged over all spills. The beam intensity was $10^{12}$ particle per spill for the run. Right: Event rate over time for the first 30 minutes of this run, clearly showing intensity spikes from the recurrent beam spills.
@@ 8 Conclusions
The commissioning of the DarkQuest EMCal was performed with test beam data. The studies on linearity, resolution, and response to various particles confirm the suitability of the EMCal and its newly developed readout electronics for the DarkQuest experiment. Light attenuation with neutral density filters and electrical attenuation were found to give similar performance. The measurements can be used as input to develop an accurate Monte Carlo simulation of the detector and readout electronics. Background rates in the real experiment were also studied and were found to be within the constraints of the existing readout system.